In a rhombus ABCD, the diagonals AC and BD intersect at the point (3,4).If the point A be (1,2),then what is the equation of the diagonal BD? 

Grade:11

1 Answers

Arun
5 years ago
In rhombus diagonals bisect at 90° degree. But they are not equal in length.
 
C will be (5, 6)
 
Now for BD-
 
slope of AC = 6-2/5-1 = 4/4 = 1
 
Hence slope of BD = -1
 
Now equation of BD 
 
y -4 = -1 (x-3)
y -4 = -x +3
x +y = 7